Population A Age (years)     Population       # Deaths          15–19              2,000 

Population A

Age (years)     Population       # Deaths         

15–19              2,000               30

20–24              4,000               50

25–29              4,000               120

Total               10,000             200

Population B

Age (years)     Population       # Deaths         

15–19              2,000               50

20–24              4,000               50

25–29              4,000               60

Total               10,000             160

  1. Which of the following is the age-adjusted rate for population B, using population A as the standard?
  1. 21 per 1,000
  2. 20 per 1,000
  3. 16 per 1,000
  4. None of these is correct.

  1. Which of the following reflects the crude mortality rate ratio for A to B?
  1. 1.05
  2. 1.25
  3. 1.89

None of these is correct
Answer #1

Answer 1-

c=(a/b)*1000

=160/10000*1000

= 16

so 16 per 1000 is the age-adjusted rate for population B

Answer 2- Total no. of cases in a given period / total no. of persons in the population

= 200/10000

=0.02 for A

for B

=160/10000

= 0.016

crude mortality rate ratio for A to B = a/b

0.02/0.016 = 1.25